OpenStudy (anonymous):

What is the equation of the line that is perpendicular to y = 2x – 4 and has a y-intercept of –2?

ganeshie8 (ganeshie8):

slope intercept form of line :- \(y = mx + b\)

ganeshie8 (ganeshie8):

\(m\) = slope \(b\) = intercept

ganeshie8 (ganeshie8):

you already have the intercept of perpendicular line : \(b = -2\) you need to find \(m\)

OpenStudy (texaschic101):

you need the negative reciprocal of the given slope

OpenStudy (anonymous):

I still dont get it :/

OpenStudy (texaschic101):

what is the slope in the equation y = 2x - 4 ?

OpenStudy (texaschic101):

In y = mx + b form, like ganshie said, the m position is the slope and the b position is the y intercept. So what is in the m position ?

OpenStudy (texaschic101):

do you know ?

OpenStudy (texaschic101):

y = mx + b (the m position is the slope) y = 2x - 4....what is in the m position ?

OpenStudy (texaschic101):

ok......the 2 is in the m position. Can you see that now ?

OpenStudy (texaschic101):

2 is the slope of the given equation. But we need a perpendicular line, so we need the negative reciprocal of that slope. All that means is " flip " the slope and change the sign. The negative reciprocal of 2, or 2/1, is -1/2 (see how I flipped the slope and changed the sign). We need to use -1/2 as our slope and -2 as our y intercept. Remember, in y = mx + b form, the m is the slope and the b is the y intercept. So the perpendicular line would be : y = -1/2x - 2

OpenStudy (anonymous):

perpendicular to y = 2x – 4 and has a y-intercept of –2: Given equation of line is 2x-y-4=0; so, equation of line perpendicular to it will be x+2y+k=0.....(I) if y-intercept is -3, then line passes through (0,-2), on putting this value of x and y in (I), k=4; therefore, equation of line is [x+2y+4=0]